# On completeness of Foliated structures, and Null Killing fields

### v1

@@ 2.4. The $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated completeness
The proof, presented in Paragraph 2.4.2, is based on the theory of Riemannian submersions, which was developed by B. O’Neil in [19] and A. Grey in [13]. We recall the main results we will need in the next subsection.
@@ 2.4.1. Prerequisites on Riemannian submersions
The horizontal distribution of a Riemannian submersion.
In all this paragraph, we consider $\pi:(E,g)\to(B,g^{\prime})$ a Riemannian submersion, and $\mathcal{H}$ the corresponding horizontal distribution. A vector $v\in T_{p}M$ is said to be horizontal if $v\in\mathcal{H}_{p}$. And a curve $\gamma:I\to E$ is said to be horizontal if for all $t\in\mathbb{R}$, $\gamma^{\prime}(t)$ is a horizontal vector, i.e. $\gamma^{\prime}(t)\in\mathcal{H}_{\gamma(t)}$.
Fact 2.7 (O’Neil, [20] Proposition 1.10).
@@
Proposition 2.8 (Corollary 1.1 page 26, [20]).
@@
Let $(M,\mathfrak{F})$ be a $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated manifold. For simplicity, we will assume that the leaves are of codimension $1$, i.e. $\dim M=n+2$ (although the proof works the same way for higher codimension). Recall that on $\mathbb{R}^{n+1}$ we have two $\mathsf{L}_{\sf{u}}(n)$-invariant objects: a vector field $V$ and a degenerate Riemannian metric $h$. Consider $V_{M}$ the pullback of $V$, it is a continuous non-vanishing vector field on $M$ tangent to $\mathfrak{F}$. Consider also $h_{M}$ the degenerate metric given by the pullback of $h$ on the subbundle $T\mathfrak{F}$. Denote by $\mathcal{V}_{M}$ the foliation tangent to $V_{M}$.
Choose an arbitrary Riemannian metric $g_{o}$ on $M$, and decompose the tangent space $TM$ orthogonaly with respect to $g_{o}$ as follows,
@@
where $Z$ is a unit vector field with respect to $g_{o}$ (it exists, up to taking a double cover).
In order to prove Theorem 2.5, we need to show (using Fact 2.2) that the developing map $D$ is a diffeomorphism from the universal cover of any leaf of $\mathfrak{F}$ onto the affine space $\mathbb{R}^{n+1}$. This will be done by constructing a particular Riemannian submersion, then we apply Subsection 2.4.1, especially, Fact 2.9.
@@
A first step toward the construction of the Riemannian submersion is to define a Riemannian metric $g$ on $M$ compatible with $h_{M}$ in the following sense
@@
for all $X,Y\in\mathcal{S}$. The metric $g$ is indeed Riemannian since the restriction of $h_{M}$ to the distribution $\mathcal{S}$ is non-degenerate and Riemannian (remember that $\mathcal{S}$ is transversal to the kernel of $h_{M}$, namely, the direction of $V_{M}$).
Fact 2.10.
Let $(M,\mathfrak{F},g)$ be a compact foliated Riemannian manifold. Then any leaf $\mathcal{F}$ with the induced Riemannian metric is geodesically complete.
@@
Let $\mathcal{F}$ be a leaf of $\mathfrak{F}$. Then $(\mathcal{F},g_{|_{\mathcal{F}}})$ is geodesically complete.
Let (~$(\widetilde{\mathcal{F}},\widetilde{g_{|_{{\mathcal{F}}_{p}}}})$ be the universal cover of $\mathcal{F}$ together with the lifted metric. The Riemannian manifold (~$(\widetilde{\mathcal{F}},\widetilde{g_{|_{{\mathcal{F}}}}})$ is geodesically complete, so we have a developing map $D:\widetilde{\mathcal{F}}\to\mathbb{R}^{n+1}$ which is a local diffeomorphism. Denote by $\mathfrak{p}$ the projection from $\mathbb{R}^{n+1}$ to the space of $\mathcal{V}$-leaves, i.e. $\mathfrak{p}:\mathbb{R}^{n+1}\to\mathbb{R}^{n+1}/\mathcal{V}\cong\mathbb{R}^{n}$. The space of $\mathcal{V}$-leaves is $\mathbb{R}^{n}$ equipped with the flat Riemannian metric, given by the projection of $h$. We denote the projected metric by $\bar{h}$. Define $\pi:=\mathfrak{p}\circ D$, this is a smooth submersion from $\widetilde{\mathcal{F}}$ to $B:=\pi(\widetilde{\mathcal{F}})$, which is an open subset of $\mathbb{R}^{n}$.
@@
The map $\pi:(\widetilde{\mathcal{F}},\widetilde{g_{|_{{\mathcal{F}}}}}))\to(B,\bar{h}_{|_{B}})$ is a Riemannian submersion.
@@
The horizontal distribution on $\widetilde{\mathcal{F}}$ is given by $\widetilde{S}$ (the lift of $S$). Let $x\in\widetilde{\mathcal{F}}$ and $y:=D(x)$, then $E_{y}:=d_{x}D(\widetilde{S}_{x})$ is transversal to the kernel of $h_{y}$ (i.e. to $\mathbb{R}V$). By construction, the metric induced on $\widetilde{S}_{x}$ by $g_{x}$ is isometric to the metric induced on $E_{y}$ by $h_{y}$. The lemma follows then from the fact that the metric induced by $h_{y}$ on any transversal is isometric to $(T_{y}\mathbb{R}^{n+1}/\mathbb{R}V,\bar{h})$. ∎
@@
We have $B=\mathbb{R}^{n}$.
@@
Let $\gamma:I\to\widetilde{\mathcal{F}}$ be a horizontal geodesic (recall that the horizontal distribution is totally geodesic by Fact 2.7). Then $\gamma$ is complete by Corollary 2.11. By Proposition 2.8, the projection of $\gamma$ by $\pi$ is a geodesic in the base space, which is an open subset of the flat euclidean space, hence $\pi\circ\gamma$ is a line segment. Since $\gamma$ is complete, this line segment is a complete affine line. This is true for any horizontal geodesic in all possible directions, hence $B=\mathbb{R}^{n}$. ∎
@@
The map $\pi:(\widetilde{\mathcal{F}},\widetilde{g_{|_{{\mathcal{F}}}}}))\to(\mathbb{R}^{n},\bar{h})$ is a trivial topological fiber bundle, i.e. we have ~$\widetilde{\mathcal{F}}\cong\mathbb{R}^{n}\times F$, with $F=\pi^{-1}(0)$.
@@
By Fact 2.9, we have that $\pi:(\widetilde{\mathcal{F}},\widetilde{g_{|_{{\mathcal{F}}}}}))\to(\mathbb{R}^{d},\bar{h})$ is a topological fiber bundle. Since the base space is contractible, the bundle is trivial (a product bundle). ∎
Theorem 2.15 (Theorem 2.5).
@@
Let $\mathcal{F}$ be a leaf of $\mathfrak{F}$, with $D:\widetilde{\mathcal{F}}\to\mathbb{R}^{n+1}$ a developing map. We will show that $D$ is a diffeomorphism. We start by the surjectivity. Recall that we have a (surjective) Riemannian submersion $\pi:\widetilde{\mathcal{F}}\to\mathbb{R}^{n}$, where $\mathbb{R}^{n}$ is the Euclidean space. Let $z\in\mathbb{R}^{n+1}$ be any point, let $x=\pi(x_{o})$ for some $x_{o}\in\widetilde{\mathcal{F}}\smallsetminus\pi^{-1}(\{z\})$. Now, consider the line segment $\gamma:=[x,\mathfrak{p}(z)]$ in $\mathbb{R}^{n}$, we set $z_{o}$ to be the other endpoint of the horizontal lift $\tilde{\gamma}$ starting at $x_{o}$. Since $\pi(z_{o})=\mathfrak{p}(z)$ we see that $D(z_{o})\in\mathcal{V}_{z}$. Moreover, the developing map sends the $\widetilde{\mathcal{V}}_{M}$ leaf containing $z_{o}$ to the $\mathcal{V}$ leaf containing $z$, i.e. $D(\widetilde{\mathcal{V}}_{M,z_{o}})\subset\mathcal{V}_{z}$. We claim that the restriction of $D$ to any $\widetilde{\mathcal{V}}_{M}$ is injective. Indeed, since the vector field $\widetilde{V}_{M}$ is the pullback of $V$ via $D$, we have that $D$ is equivariant with respect to their flows, i.e. for $\tilde{p}\in\widetilde{F}$ we have
$D(\phi_{\widetilde{V}_{M}}^{t}(\tilde{p}))=\phi_{V}^{t}(D(\tilde{p})).$
Moreover, the flow of $V$ acts freely on the affine space, in particular on its orbits (affine lines). The claim follows. Now, since $M$ is compact the flow of $\widetilde{V}_{M}$ is complete, so, the restriction of developing map to a $\widetilde{\mathcal{V}}_{M}$ leaf is also surjective. Hence, there exists some $z^{\prime}\in\widetilde{\mathcal{V}}_{M,z_{o}}$ such that $D(z^{\prime})=z$.
For the injectivity of $D$, assume that there are two distinct points $x_{o},y_{o}\in\widetilde{\mathcal{F}}$ (necessarily in different $\widetilde{\mathcal{V}}_{M}$ leaves) such that $D(x_{o})=D(y_{o})$. Hence, $\pi(x_{o})=\pi(y_{o})=x$, which means that the fiber of $x$ contains at least two $\widetilde{\mathcal{V}}_{M}$ leaves. On other hand, Corollary 2.14 together with the fact that $\widetilde{\mathcal{F}}$ is connected implies that the fiber is necessarily connected. Therefore contradiction. ∎
Corollary 2.16.
Let $(M,\mathfrak{F})$ be a compact $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated manifold. Then the leaves of $\mathfrak{F}$ are geodesically complete with respect to the flat affine connection.
Proof.
On a leaf $\mathcal{F}$ of $\mathfrak{F}$ we can define the “pullback” of the flat affine connection (since it is invariant by $\mathsf{L}_{\sf{u}}(n)$). However, since $\mathcal{F}$ is $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-complete we deduce by Fact 2.2 that it is geodesically complete. ∎
@@
Recall that the invariant objects for this structure are the $1$-form $\omega:=dx_{n+1}$ and the codimension $1$ foliation $\mathfrak{G}$ tangent to the kernel of this form. Denote by $\mathfrak{G}_{M}$ the foliation on $M$, tangent to $\mathfrak{F}$, given by the pullback of $\mathfrak{G}$. And denote by $\omega_{M}$ the $1$-form on $M$ given by the pullback of $\omega$. Similarly to the proof of Theorem 2.5, we show that the leaves of $\mathfrak{G}_{M}$ are complete. On the other hand, there is a vector field $W$ of $M$ such that $\omega_{M}(W)=1$. Indeed, $W$ can be chosen as a global section of a tangent line bundle, this global section exists since the fibers are contractible. Consider the $1$-from $\widetilde{\omega}_{M}$ and the vector field $\widetilde{W}$ on $\widetilde{M}$. Let $\tilde{p}\in\widetilde{M}$. Since $\omega$ is $\mathsf{L}_{\sf{u}}^{*}(n)$-invariant, we have the following formula
$D\circ\phi_{\widetilde{W}}^{t}(\widetilde{\mathcal{G}}_{M}(\tilde{p}))=\phi_{\partial x_{n+1}}^{t}\circ D(\widetilde{\mathcal{G}}_{M}(\tilde{p}))=\phi_{\partial x_{n+1}}^{t}(\mathcal{G}(D(\tilde{p}))),$
where $\phi_{\widetilde{W}}^{t}$ is the flow of $\widetilde{W}$, where $\mathcal{G}(p)$ (resp. $\mathcal{G}_{M}(p)$) is the leaf of $\mathfrak{G}$ (resp. $\mathfrak{G}_{M}$) at $p$. From this we see that $D$ is surjective since $\widetilde{W}$ is complete (by compactness of $M$), and injective since the $\widetilde{W}$-flow acts transitively on the space of $\widetilde{\mathfrak{G}}_{M}$-leaves. ∎
2.5. Application to compact pp-waves

### v2

@@ 2.4. $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated completeness
The proof, presented in Paragraph 2.4.2, is based on the theory of Riemannian submersions, developed by B. O’Neil in [29] and A. Grey in [16]. In the next paragraph, we will recall the main results needed for the proof.
@@ 2.4.1. Prerequisites on Riemannian submersions
Consider $\pi:(E,g)\to(B,g^{\prime})$ a Riemannian submersion, and $\mathcal{H}$ the corresponding horizontal distribution. A vector $v\in T_{p}M$ is said to be horizontal if $v\in\mathcal{H}_{p}$. And a curve $\gamma:I\to E$ is said to be horizontal if for all $t\in\mathbb{R}$, $\gamma^{\prime}(t)$ is a horizontal vector, i.e. $\gamma^{\prime}(t)\in\mathcal{H}_{\gamma(t)}$.
We now state three facts that will be used throughout the section.
Fact 2.12 (O’Neil, [31] Proposition 1.10).
@@
Fact 2.13 (Corollary 1.1 page 26, [31]).
@@
In order to prove Theorem 2.10, we need to show that the developing map ${\mathsf{Dev}}$ is a diffeomorphism from the universal cover of any leaf of $\mathfrak{F}$ onto the affine space $\mathbb{R}^{n+1}$. This will be done by constructing a particular Riemannian submersion, and applying Subsection 2.4.1.
Let $(M,\mathfrak{F})$ be a $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated manifold. For simplicity, we will assume that the leaves are of codimension $1$, i.e. $\dim M=n+2$ (although the proof works the same way for higher codimension). Recall that on $\mathbb{R}^{n+1}$ we have two $\mathsf{L}_{\sf{u}}(n)$-invariant objects: a vector field $V$ and a degenerate Riemannian metric $h$. Consider $V_{M}$ the pullback of $V$, it is a continuous non-vanishing vector field on $M$ tangent to $\mathfrak{F}$. Denote by $\mathcal{V}_{M}$ the foliation tangent to $V_{M}$. Consider also $h_{M}$ the degenerate metric given by the pullback of $h$ on the subbundle $T\mathfrak{F}$.
Choose an arbitrary Riemannian metric $g_{0}$ on $M$, and decompose the tangent bundle $TM$ orthogonally with respect to $g_{0}$ as follows,
@@
where $\mathcal{S}$ is an $n$-dimensional distribution tangent to $\mathfrak{F}$ and transversal to $V_{M}$, and $Z$ is a $g_{0}$-unit vector field (it exists, up to taking a double cover).
@@
We define a Riemannian metric $g$ on $M$, compatible with $h_{M}$ in the following sense
@@
for all $X,Y\in\mathcal{S}$. The metric $g$ is indeed Riemannian, since the restriction of $h_{M}$ to the distribution $\mathcal{S}$ is non-degenerate and Riemannian (recall that $\mathcal{S}$ is transversal to $V_{M}$, which spans the radical of $h_{M}$).
Fact 2.15.
Let $(M,\mathfrak{F},g)$ be a compact foliated Riemannian manifold. Then any leaf $F$ with the induced Riemannian metric is geodesically complete.
@@
Let $F$ be a leaf of $\mathfrak{F}$. Then $(F,g_{|_{F}})$ is geodesically complete.
Proof.
This follows from Fact 2.15. ∎
Let $(\widetilde{F},\widetilde{g_{|_{{F}}}})$ be the universal cover of $F$, endowed with the lifted Riemannian metric. The Riemannian manifold $(\widetilde{F},\widetilde{g_{|_{{F}}}})$ is geodesically complete. Let ${\mathsf{Dev}}:\widetilde{F}\to\mathbb{R}^{n+1}$ be a developing map for the leaf $F$. Let $\mathfrak{p}:\mathbb{R}^{n+1}\to\mathbb{R}^{n+1}/\mathcal{V}\cong\mathbb{R}^{n}$ be the projection onto the space of $\mathcal{V}$-leaves. The latter is identified to $\mathbb{R}^{n}$, equipped with the flat Riemannian metric given by the projection of $h$ by $\mathfrak{p}$. We denote the projected metric by $\bar{h}$. Define $\pi:=\mathfrak{p}\circ{\mathsf{Dev}}$, this is a smooth submersion from $\widetilde{F}$ to $B:=\pi(\widetilde{F})\subset\mathbb{R}^{n}$, an open subset of $\mathbb{R}^{n}$.
@@
The map $\pi:(\widetilde{F},\widetilde{g_{|_{{F}}}}))\to(B,\bar{h}_{|_{B}})$ is a Riemannian submersion.
@@
The horizontal distribution on $\widetilde{F}$ is given by $\widetilde{\mathcal{S}}$, the lift of $\mathcal{S}_{|F}$ to $\widetilde{F}$. Let $x\in\widetilde{F}$ and let $y:={\mathsf{Dev}}(x)$. Then, $E_{y}:=d_{x}{\mathsf{Dev}}(\widetilde{\mathcal{S}}_{x})$ is a vector subspace of $T_{y}\mathbb{R}^{n+1}$ transversal to the radical of $h_{y}$ (i.e. to $\mathbb{R}V$). By construction, the metric induced on $\widetilde{\mathcal{S}}_{x}$ by $g$ is isometric to the metric induced on $E_{y}$ by $h$. The lemma follows then from the fact that the metric induced by $h$ on any vector subspace of $T_{y}\mathbb{R}^{n+1}$ transversal to $V$ is isometric to $(T_{y}\mathbb{R}^{n+1}/\mathbb{R}V,\bar{h}_{\mathfrak{p}(y)})$. ∎
@@
$(B,\bar{h}_{|_{B}})$ is complete, in particular, we have $B=\mathbb{R}^{n}$.
@@
Let $\gamma:I\to\widetilde{F}$ be a horizontal geodesic (recall that the horizontal distribution is totally geodesic by Fact 2.12). Then $\gamma$ is complete by Corollary 2.16. Moreover, by Fact 2.13, the projection of $\gamma$ by $\pi$ is a geodesic in the base space, which is an open subset of the flat euclidean space, hence $\pi\circ\gamma$ is an affine segment. Since $\gamma$ is complete, this affine segment is a complete affine line. This is true for any horizontal geodesic in all possible directions, hence $B=\mathbb{R}^{n}$. ∎
@@
The map $\pi:(\widetilde{F},\widetilde{g_{|_{{F}}}}))\to(\mathbb{R}^{n},\bar{h})$ is a trivial topological fiber bundle, i.e. we have $\widetilde{F}\cong\mathbb{R}^{n}\times l$, with $l=\pi^{-1}(0)$.
@@
By Fact 2.14, we have that $\pi:(\widetilde{F},\widetilde{g_{|_{{F}}}}))\to(\mathbb{R}^{n},\bar{h})$ is a topological fiber bundle. Since the base space is contractible, the bundle is trivial (a product bundle). ∎
We are now able to prove the main theorem of this section, Theorem 2.10.
Proof of Theorem 2.10.
Let $F$ be a leaf of $\mathfrak{F}$, with ${\mathsf{Dev}}:\widetilde{F}\to\mathbb{R}^{n+1}$ a developing map. We will show that ${\mathsf{Dev}}$ is a diffeomorphism onto $\mathbb{R}^{n+1}$. We start by the surjectivity. Recall that the Riemannian submersion $\pi=\mathfrak{p}\circ{\mathsf{Dev}}:\widetilde{F}\to\mathbb{R}^{n}$, where $\mathbb{R}^{n}$ is the Euclidean space, is surjective. Let $z\in\mathbb{R}^{n+1}$ be any point, and let $x_{o}\in\widetilde{F}\smallsetminus\pi^{-1}(\{\mathfrak{p}(z)\})$. Set $x:=\pi(x_{o})$. Consider the line segment $\gamma:=[x,\mathfrak{p}(z)]$ in $\mathbb{R}^{n}$, and let $\tilde{\gamma}$ be the horizontal lift of $\gamma$ by $\pi$, starting at $x_{o}$. Denote by $z_{o}$ the other endpoint of $\tilde{\gamma}$ in $\widetilde{F}$. Since $\pi(z_{o})=\mathfrak{p}(z)$, we see that ${\mathsf{Dev}}(z_{o})\in\mathcal{V}(z)$, where $\mathcal{V}(z)$ denotes the $\mathcal{V}$-leaf containing $z$. Moreover, the developing map sends the $\widetilde{\mathcal{V}}_{M}$-leaf containing $z_{o}$ to the $\mathcal{V}$-leaf containing $z$, i.e. ${\mathsf{Dev}}(\widetilde{\mathcal{V}}_{M}(z_{o}))\subset\mathcal{V}(z)$. We claim that the restriction of ${\mathsf{Dev}}$ to any $\widetilde{\mathcal{V}}_{M}$-leaf is injective. Indeed, since the vector field $\widetilde{V}_{M}$ is the pullback of $V$ via ${\mathsf{Dev}}$, it follows that ${\mathsf{Dev}}$ is equivariant with respect to their respective flows, i.e. for any $\tilde{p}\in\widetilde{F}$ we have
${\mathsf{Dev}}(\phi_{\widetilde{V}_{M}}^{t}(\tilde{p}))=\phi_{V}^{t}({\mathsf{Dev}}(\tilde{p})).$
And now, since the flow of $V$ acts freely (by translations) on the affine space $\mathbb{R}^{n+1}$, and in particular, on its orbits (affine lines), the claim follows. On the other hand, since $M$ is compact, the flow of $\widetilde{V}_{M}$ is complete. Therefore, by the equivariance formula, the restriction of the developing map to any $\widetilde{\mathcal{V}}_{M}$-leaf is also surjective. The surjectivity of ${\mathsf{Dev}}$ follows.
For the injectivity, assume that there are two distinct points $x_{o},y_{o}\in\widetilde{F}$ (necessarily in different $\widetilde{\mathcal{V}}_{M}$-leaves) such that ${\mathsf{Dev}}(x_{o})={\mathsf{Dev}}(y_{o})$. Hence, $\pi(x_{o})=\pi(y_{o})=x$, which means that the fiber of $x$ contains at least two $\widetilde{\mathcal{V}}_{M}$-leaves. However, Corollary 2.19 together with the fact that $\widetilde{F}$ is connected imply that the fiber is necessarily connected, leading to a contradiction. ∎
Given a $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated manifold $(M,\mathfrak{F})$, the pullback to $T\mathfrak{F}$ of the flat affine connection of $\mathbb{R}^{n+1}$ defines a flat affine connection on the leaves of $\mathfrak{F}$. And we have
Corollary 2.20.
Let $(M,\mathfrak{F})$ be a compact $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated manifold. Then the leaves of $\mathfrak{F}$ are geodesically complete with respect to the induced flat affine connection.
@@
This is a consequence of Theorem 2.10 and Fact 2.2. ∎
For the dual geometry, we need the following fact:
Fact 2.21 (Proposition $8$ [23]).
Let $M$ be a compact manifold endowed with a nowhere vanishing closed $1$-form $\omega$. Let $\widetilde{\omega}$ be the pullback of $\omega$ to the universal cover $\widetilde{M}$, and $Z$ a vector field on $\widetilde{M}$ such that $\widetilde{\omega}(Z)=1$. Then $\widetilde{M}$ is diffeomorphic to ×$\mathbb{R}\times\widetilde{F}$ where $\widetilde{F}$ is a leaf of the foliation $\mathfrak{F}_{\widetilde{M}}$ tangent to $\ker(\widetilde{\omega})$. And the flow of $Z$ acts simply transitively on the space of leaves, identified to $\mathbb{R}$, and maps a leaf to another diffeomorphically.
@@
In the basis $(e_{1},\ldots,e_{n+1})$ of $\mathbb{R}^{n+1}$ where the matrix group $\mathsf{L}_{\sf{u}}^{*}(n)$ is represented, the $1$-form $\omega:=e_{n+1}^{*}$ as well as the codimension $1$ foliation $\mathfrak{G}$ tangent to $\ker\omega$ are $\mathsf{L}_{\sf{u}}^{*}(n)$-invariant. Denote by $\mathfrak{G}_{M}$ the foliation on $M$, tangent to $\mathfrak{F}$, given by the pullback of $\mathfrak{G}$. And denote by $\omega_{M}$ the $1$-form on $M$ given by the pullback of $\omega$. There exists a vector field $W$ on $M$ such that $\omega_{M}(W)=1$. In fact, $W$ can be chosen as a global section of a tangent affine subbundle with contractible fibers, ensuring the existence of the section. We denote by $\widetilde{\omega}_{M}$, $\widetilde{W}$, $\widetilde{\mathfrak{G}_{M}}$ the lifts of the corresponding objects to the universal cover $\widetilde{M}$. We denote the flow of a vector field $U$ by $\phi_{U}^{t}$. Since $M$ is compact, $W$ is complete. Fix a leaf $F$ of $\mathfrak{F}$, and consider a connected component $\widetilde{F}$ of its lift to $\widetilde{M}$. We have a developing map ${\mathsf{Dev}}:\widetilde{F}\to\mathbb{R}^{n+1}$. By Fact 2.21, the flow of $\widetilde{W}$, restricted to $\widetilde{F}$, acts simply transitively on the space of $\widetilde{\mathfrak{G}_{M}}_{|\widetilde{F}}$-leaves, and maps one leaf diffeomorphically to another. And since $\omega$ is $\mathsf{L}_{\sf{u}}^{*}(n)$-invariant, we have the following equivariance formula
${\mathsf{Dev}}\circ\phi_{\widetilde{W}}^{t}=\phi_{e_{n+1}}^{t}\circ{\mathsf{Dev}}.$
On the other hand, we can show similarly to the proof of Theorem 2.10 that ${\mathsf{Dev}}$, restricted to a leaf of $\widetilde{\mathfrak{G}_{M}}$ through $p\in\widetilde{F}$, is bijective onto the leaf of $\mathfrak{G}$ through ${\mathsf{Dev}}(p)\in\mathbb{R}^{n+1}$. Fix a leaf of $\widetilde{\mathfrak{G}_{M}}$ in $\widetilde{F}$. The equivariance formula applied to this leaf shows that the developing map is surjective, thus, bijective on all $\widetilde{F}$, which ends the proof. ∎
2.5. Applications
In what follows, we discuss applications of Theorem 2.10.
2.5.1. Compact pp-waves.